# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ 15.2 Geometric Quantization via Degenerate Lefschetz Thimbles
Because the integration variables are coupled complex holomorphic twistors, the vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy). This complex holomorphic geometry natively upgrades the standard square-root branch cut into a strictly simple pole, precisely generating the discrete S-matrix mass poles of the meson spectrum without ad hoc approximations.
@@ 15.3 Catastrophes, flat valleys, and topological barriers
To dynamically generate an exact simple mass pole $\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat zero-mode that extends infinitely far in configuration space. This exact infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields as the boundary parameter winds around the enclosed twistor poles.
@@
The integrity of this topological winding number is rigorously protected by an infinite topological action barrier. If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside. This creates a severe pinch singularity, strictly trapping the integration contour and locking the path integral into a conserved topological sector.
15.4 Poles inside the disk and the monodromies
For a saddle point with a single pole located strictly inside the unit disk ($|z_{0}|<1$), the total effective action accumulates strictly additively over multiple windings $k$. The Liouville bulk action simply duplicates, and the logarithmic boundary primitive picks up a strict topological phase shift at every cycle. Together, these terms yield an infinite geometric series $\sum\exp\left(\mathrm{i}k\Delta S_{eff}\right)$. The macroscopic divergence of this series generates the simple meson pole precisely when the exact Bohr-Sommerfeld quantization condition $\Delta S_{eff}=2\pi n$ is satisfied.
15.5 The Exact One-Pole Solution and Minkowski Transition
For the fundamental meson trajectory, we consider in detail the specific case of a single pole. To implement the twisted boundary condition for $\lambda(z)$ while freezing $\mu(z)$ to enforce the symmetrization over SD and ASD surfaces, we utilize the exact chiral 1-pole configuration with fractional power $a$, shifted to a pole position $z_{0}$:
$\lambda(z)=R(z-z_{0})^{-a}\begin{pmatrix}1\\ 0\end{pmatrix},\qquad\mu(z)=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 1\end{pmatrix}$ (15.2)
As we trace the boundary loop $z\to z\exp\left(2\pi\mathrm{i}\right)$, the active twistor acquires the fractional monodromy $\lambda\to\exp\left(-2\pi\mathrm{i}a\right)\lambda$. Equating this to the physical spin-1/2 transformation $\lambda\to\exp\left(-\mathrm{i}\phi/2\right)\lambda$, we establish the exact geometric identification $\phi=4\pi a$. The angular momentum projection operator thereby injects the canonical phase $\exp\left(4\pi\mathrm{i}Ja\right)$.
To evaluate the bulk action properly, we must transition to the physical Minkowski metric on the symmetrized chiral surface, as established in D. We switch to the Minkowski metric on the surface parameterized by $z=\exp\left(\mathrm{i}\sigma_{M}+\tau_{E}\right)$, describing the Nambu-Goto metric $g_{ab}$ via the coordinate transformation $\exp\left(2\rho_{E}\right)d^{2}z=R^{2}\frac{d^{2}z}{(z\bar{z})^{a}}=R^{2}\mathrm{i}d\sigma_{M}d\tau_{E}$. Upon double Wick rotation and symmetrization (taking the parity-even arithmetic mean of the Left and Right chiral configurations), the conformal factor of the twisted metric dynamically restricts the causal domain. Using the modified radial coordinate $\tilde{\sigma}_{M}=(1-a)\sigma_{M}$, the physical speed-of-light boundary occurs exactly where this conformal metric vanishes, at $\tilde{\sigma}_{M}=\pi/2$. The proper Minkowski Area integral factorizes in the light-cone coordinates $z,\bar{z}$:
$\text{Area}(a)$ $=\frac{\mathrm{i}R^{2}}{2}\int_{z\overline{z}<1}dz\wedge d\overline{z}(z\overline{z})^{-a}$
$=\frac{-\mathrm{i}R^{2}}{2}\oint\frac{dz}{z}\int_{0}^{1}d(z\overline{z})(z\overline{z})^{-a}$
$=\frac{-\mathrm{i}R^{2}}{2(1-a)}\oint dzz^{-1}=\frac{\pi R^{2}}{1-a}$ (15.3)
Multiplying by the physical string tension $\sigma$, the macroscopic Area action evaluates exactly to:
$S_{M,Area}=\frac{\pi\sigma R^{2}}{1-a}$ (15.4)
Simultaneously, the Liouville action evaluates the kinetic gradient term. Crucially, while the target space metric stretches with the twist, the topological Liouville anomaly resides strictly on the rigidly bounded fundamental domain of the reference worldsheet, $\sigma_{M}\in[0,\pi/2]$:
$\int|\nabla\rho|^{2}d^{2}z$ $=a^{2}\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}d\sigma_{M}$
$=a^{2}(2\pi)\left(\frac{\pi}{2}\right)=\pi^{2}a^{2}$ (15.5)
Multiplying by the universal anomaly coefficient $\frac{1}{12\pi}$, the Minkowski kinetic action yields the strictly positive, rigid Lüscher shift:
$S_{M,kin}=+\frac{\pi a^{2}}{12}$ (15.6)

### 23 Mar 2026 v4

@@ 15.2 Geometric Quantization via Degenerate Lefschetz Thimbles
Because the integration variables are coupled complex holomorphic twistors, the vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of two real directions on the Lefschetz thimble (a strict corank-2 degeneracy). This complex holomorphic geometry natively upgrades the standard square-root branch cut into a simple pole, precisely generating the discrete S-matrix mass poles of the meson spectrum without ad hoc approximations.
@@ 15.3 Catastrophes, flat valleys, and topological barriers
To dynamically generate a simple mass pole $\int d^{2}w\exp\left(-\bar{w}Q(E)w\right)\propto 1/Q(E)\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat complex zero-mode $w$ that extends infinitely far in configuration space. This infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields as the boundary parameter winds around the enclosed twistor poles.
@@
The integrity of this topological winding number is protected by an infinite topological action barrier. If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside. This creates a pinch singularity, trapping the integration contour and locking the path integral into a conserved topological sector. We demonstrate that in D on the example of a single pole in twistor field.
16 Minkowski Minimal Surface and Meson Spectrum
To implement this flat valley scenario and compute the spectrum of massive mesons, we construct the minimal surface directly in four-dimensional Minkowski space $\mathbb{R}^{1,3}$. This is more subtle than just a straightforward analytic continuation of the twistor parametrization of a Euclidean minimal surface.
16.1 Additive Factorization in Minkowski Space
To construct a surface in Minkowski space, we utilize the generalized Weierstrass representation of Konopelchenko and Landolfi (K&L Theorem 5.1) [16]. The target-space coordinate differential, represented as the Hermitian matrix $d\mathbf{X}=dX^{0}I+dX^{1}\sigma_{1}+dX^{2}\sigma_{2}+dX^{3}\sigma_{3}$, is parameterized by an additive sum of two independent, 2-component complex spinors:
$d\mathbf{X}=\phi(\xi)\phi^{\dagger}(\xi)d\xi+\psi(\eta)\psi^{\dagger}(\eta)d\eta.$ (16.1)
Here, the worldsheet light-cone coordinates are defined as $\xi=\tau+\theta$ and $\eta=\tau-\theta$, where $\tau$ is the periodic time on the fundamental torus and $\theta$ is the spatial radial parameter. Because $d\mathbf{X}$ is constructed as the sum of rank-1 Hermitian matrices, the target-space coordinates $X^{\mu}$ are real.
We use the $\psi,\phi$ notations instead of Euclidean $\lambda,\mu$. The dependence of the left/right spinors $\psi/\phi$ on a single coordinate $\tau\pm\theta$ is the Minkowski analog of Euclidean holomorphicity. The classical minimal surface equations in Minkowski reduce to d’Alembert equations $\partial_{+}\partial_{-}X=0$ combined with Virasoro constraints $\det d\mathbf{X}=0$, equivalent to the null vector condition $dX_{\mu}^{2}=0$ in Minkowski space. Both of these conditions are satisfied in the above twistor representation.
16.2 Complex Eigenvectors and Target-Space Monodromy
To incorporate the rotation of the meson, we require the coordinate matrix $d\mathbf{X}$ to be unitary transformed by a spatial rotation matrix when the time coordinate winds around its fundamental cycle on the torus ($\tau\to\tau+2\pi$).
Because the left-moving ($\xi$) and right-moving ($\eta$) dependencies are decoupled, we achieve this target-space monodromy by defining the component spinors $\phi$ and $\psi$ as complex eigenvectors of the $SU(2)$ rotation group. To center the string at the origin, we introduce a constant $\pi/4$ phase shift and scale the spinors by the macroscopic geometric scale $R$:
$\phi(\xi)=\sqrt{\frac{R}{2}}\begin{pmatrix}\exp\left(i(a\xi+\pi/4)\right)\\ \exp\left(-i(a\xi+\pi/4)\right)\end{pmatrix}$,
$\psi(\eta)=\sqrt{\frac{R}{2}}\begin{pmatrix}\exp\left(i(a\eta-\pi/4)\right)\\ \exp\left(-i(a\eta-\pi/4)\right)\end{pmatrix}$. (16.2)
Under arbitrary temporal shift $\tau\to\tau+\Delta\tau$, the arguments shift as $\xi\to\xi+\Delta\tau$ and $\eta\to\eta+\Delta\tau$. The column vectors are left-multiplied by the diagonal rotation matrix:
$\phi(\tau+\Delta\tau)=M(2a\Delta\tau)\phi(\tau),$
$\psi(\eta+\Delta\tau)=M(2a\Delta\tau)\psi(\eta)$
$\text{where }M(\alpha)=\begin{pmatrix}\exp\left(\mathrm{i}\alpha/2\right)&0\\ 0&\exp\left(-\mathrm{i}\alpha/2\right)\end{pmatrix}$. (16.3)
Substituting this into the additive factorization, the target-space coordinate differential rotates as:
$d\mathbf{X}(\tau+\Delta\tau)=M(2a\Delta\tau)\,d\mathbf{X}(\tau)\,M^{\dagger}(2a\Delta\tau).$ (16.4)
Note that this applies to arbitrary temporal shifts $\Delta\tau$, not just the fundamental period $2\pi$. Later, we discuss the implications of this extended symmetry. This twisted boundary condition will allow us to project the quark loop amplitude to a fixed angular momentum $J$ by means of the Fourier integral (15.1) with $\alpha=4\pi a$.
16.3 Real Coordinates and 2D Rotation
Substituting the complex eigenvectors into the coordinate differentials, we evaluate the diagonal components:
$dX^{0}+dX^{3}$ $=|\phi_{1}|^{2}d\xi+|\psi_{1}|^{2}d\eta=R\,d\tau.$ (16.5)
$dX^{0}-dX^{3}$ $=|\phi_{2}|^{2}d\xi+|\psi_{2}|^{2}d\eta=R\,d\tau.$ (16.6)
This yields $dX^{0}=R\,d\tau$ and $dX^{3}=0$. The center of mass is stationary. Evaluating the off-diagonal components yields the transverse spatial coordinates:
$dX^{1}-idX^{2}=\phi_{1}\bar{\phi}_{2}d\xi+\psi_{1}\bar{\psi}_{2}d\eta$
$=\frac{R}{2}\exp\left(i(2a\xi+\pi/2)\right)d\xi+\frac{R}{2}\exp\left(i(2a\eta-\pi/2)\right)d\eta$
$=\frac{iR}{2}\exp\left(2ia\xi\right)d\xi-\frac{iR}{2}\exp\left(2ia\eta\right)d\eta.$ (16.7)
Integrating this differential produces:
$X^{1}-iX^{2}=\frac{R}{4a}(\exp\left(2ia\xi\right)-\exp\left(2ia\eta\right))$
$=\frac{R}{4a}\exp\left(2ia\tau\right)(\exp\left(2ia\theta\right)-\exp\left(-2ia\theta\right))$
$=\frac{iR}{2a}\exp\left(2ia\tau\right)\sin(2a\theta).$ (16.8)
Separating the real and imaginary parts yields real spatial coordinates $X^{1}$ and $X^{2}$. The spatial parameter $\theta$ acts as a signed radial coordinate originating from the stationary center of mass at $\theta=0$. The symmetric interval $\theta\in[-\theta_{b},\theta_{b}]$ parameterizes the open string from the antiquark at $\theta_{b}$, straight through the center of mass, to the quark at $\theta_{b}$, precisely once without duplication. The string is a straight line spanning the transverse plane, sweeping a 2D circular rotation with a twist of $\alpha=4\pi a$.
16.4 The Vanishing Metric and Spatial Boundaries
The induced metric on the worldsheet is determined by the determinant of the target-space coordinate differentials. In this formulation, the conformal factor evaluates to the absolute square of the cross-term determinant:
$\Omega^{2}=|\phi_{1}\psi_{2}-\phi_{2}\psi_{1}|^{2}.$ (16.9)
Evaluating this for the given complex eigenvectors:
$\phi_{1}\psi_{2}-\phi_{2}\psi_{1}$
$=\frac{R}{2}(\exp\left(ia(\xi-\eta)+i\pi/2\right)-\exp\left(-ia(\xi-\eta)-i\pi/2\right))$
$=\frac{R}{2}(i\exp\left(2ia\theta\right)+i\exp\left(-2ia\theta\right))=iR\cos(2a\theta).$ (16.10)
Taking the absolute square $\Omega^{2}=R^{2}\cos^{2}(2a\theta)$, the induced metric on the worldsheet is:
$ds^{2}=R^{2}\cos^{2}(2a\theta)(d\tau^{2}-d\theta^{2}).$ (16.11)
In Minkowski space, the Lorentzian signature $(d\tau^{2}-d\theta^{2})$ is maintained.
This metric vanishes at the roots $\theta=\pm\frac{\pi}{4a}$, which correspond to the string endpoints moving at the speed of light ($ds^{2}=0$). For massive endpoint quarks ($m_{q}>0$), the physical string is truncated at boundaries $\pm\theta_{b}$ inside this null radius ($0\leq\theta_{b}<\frac{\pi}{4a}$), ensuring a positive proper time.
16.5 Classical Action Components over the Torus
We evaluate the classical WKB action components integrated over the fundamental torus temporal period $\tau\in[0,2\pi]$ and the bounded spatial domain $\theta\in[-\theta_{b},\theta_{b}]$. We define the $a$-independent boundary phase parameter $\beta=2a\theta_{b}$.
1. The Target Time Boundary Term ($E\Delta X^{0}$):
The canonical energy $E$ conjugates to the total center-of-mass target time advance $\Delta X^{0}$. Integrating this over the temporal period gives:
$E\Delta X^{0}=E\int_{0}^{2\pi}R\,d\tau=2\pi ER.$ (16.12)
2. The Minimal Area Term ($S_{\text{Area}}$):
The area of the minimal surface integrates the conformal factor over the dynamically bounded spatial domain:
$S_{\text{Area}}$ $=\sigma R^{2}\int_{0}^{2\pi}d\tau\int_{-\theta_{b}}^{\theta_{b}}\cos^{2}(2a\theta)\,d\theta$
$=2\pi\sigma R^{2}\int_{-\theta_{b}}^{\theta_{b}}\frac{1+\cos(4a\theta)}{2}\,d\theta$
$=\frac{\pi\sigma R^{2}}{a}\left(\beta+\frac{1}{2}\sin(2\beta)\right).$ (16.13)
3. The Proper Mass Integral ($S_{\text{mass}}$):
The constituent quarks of bare mass $m_{q}$ propagate along the two boundaries $\pm\theta_{b}$. Evaluating the proper time $ds=R\cos(\beta)d\tau$ accumulated over the torus period for both endpoints yields:
$S_{\text{mass}}=2m_{q}\int_{0}^{2\pi}R\cos(\beta)\,d\tau=4\pi m_{q}R\cos(\beta).$ (16.14)
4. The Liouville term:
To complete the evaluation of the geometric phase, we compute the contribution from the Liouville anomaly on the minimal surface. In the conformal gauge, the Liouville action is given by the worldsheet integral
$S_{\rm Liouv}=\frac{1}{12\pi}\int d\tau d\theta\left((\partial_{\theta}\rho)^{2}-(\partial_{\tau}\rho)^{2}\right).$ (16.15)
Using the conformal metric scale factor induced by our twistor parameterization in Minkowski space, $e^{2\rho}=R^{2}\cos^{2}(2a\theta)$, the Liouville field is purely spatial, $\rho(\theta)=\ln R+\ln\cos(2a\theta)$. Thus, the proper time derivative $\partial_{\tau}\rho$ vanishes, and the spatial gradient squared evaluates to $(\partial_{\theta}\rho)^{2}=4a^{2}\tan^{2}(2a\theta)$.
Integrating this over one full period of the proper time ($\Delta\tau=2\pi$) and symmetrically across the spatial boundaries of the string ($\theta\in[-\theta_{b},\theta_{b}]$), and accounting for the overall minus sign of the measure, we obtain the anomaly defect:
$\Delta S_{\rm Liouv}$ $=-\frac{1}{12\pi}(2\pi)\int_{-\theta_{b}}^{\theta_{b}}4a^{2}\tan^{2}(2a\theta)\,d\theta$
$=-\frac{2a^{2}}{3}\int_{-\theta_{b}}^{\theta_{b}}\left(\sec^{2}(2a\theta)-1\right)d\theta.$ (16.16)
Evaluating this integral and substituting the dynamical boundary velocity parameter $\beta=2a\theta_{b}$ yields:
$\Delta S_{\rm Liouv}=-\frac{2a}{3}(\tan\beta-\beta).$ (16.17)
16.6 Path-Ordered Dirac Trace and Spin Holonomy
To relate the continuous twistor geometry to the discrete meson spin states, we compute the path-ordered product of the Dirac projectors along the spatial boundary of the quark loop. Following the staggered spinor diagonalization of Section 12.6, the fundamental link matrix evaluated in Minkowski space is $\mathcal{M}(\tau)=i\gamma_{\mu}v^{\mu}(\tau)$, where $v^{\mu}=\partial_{\tau}X^{\mu}$ is the target-space velocity.
Substituting the twistor coordinate differentials yields the velocity components at the boundary $\theta_{b}$:
$v^{0}=R,\qquad v^{3}=0,$
$v^{1}=-R\sin(\beta)\cos(2a\tau),$
$v^{2}=R\sin(\beta)\sin(2a\tau).$ (16.18)
Defining the transverse velocity fraction $S=\sin(\beta)$, the invariant squared velocity is $v_{\mu}v^{\mu}=R^{2}(1-S^{2})=R^{2}\cos^{2}(\beta)$. In the chiral Weyl representation, the Dirac matrix projection takes the off-diagonal form $\mathcal{M}(\tau)=i\begin{pmatrix}0&\bar{V}(\tau)\\ V(\tau)&0\end{pmatrix}$, with the $2\times 2$ blocks defined as:
$V(\tau)=R\begin{pmatrix}1&-S\exp\left(2ia\tau\right)\\ -S\exp\left(-2ia\tau\right)&1\end{pmatrix}$,
$\bar{V}(\tau)=R\begin{pmatrix}1&S\exp\left(2ia\tau\right)\\ S\exp\left(-2ia\tau\right)&1\end{pmatrix}$. (16.19)
Applying the staggered transformation $\Omega_{n}$ alternating between $I_{4}$ and $\gamma_{0}$, the product of two adjacent links at step $j$ assumes the block-diagonal form:
$\tilde{\mathcal{M}}_{j+1}\tilde{\mathcal{M}}_{j}=-\begin{pmatrix}\bar{V}(\tau_{j+1})V(\tau_{j})&0\\ 0&V(\tau_{j+1})\bar{V}(\tau_{j})\end{pmatrix}$. (16.20)
Factoring out the scalar invariant $v^{2}$, the continuous limit of the path-ordered exponential separates into two independent $2\times 2$ evolution operators, $W(\tau)$ and $\bar{W}(\tau)$. For the upper block, expanding $\bar{V}(\tau+d\tau)V(\tau)$ yields the evolution equation $W^{\prime}(\tau)=A(\tau)W(\tau)$, governed by the connection matrix:
$A(\tau)=\frac{2iaS}{1-S^{2}}\begin{pmatrix}-S&\exp\left(2ia\tau\right)\\ -\exp\left(-2ia\tau\right)&S\end{pmatrix}$. (16.21)
This system is solved by transforming to a co-rotating frame. Defining $W(\tau)=U(\tau)\tilde{W}(\tau)$ with $U(\tau)=\text{diag}(\exp\left(ia\tau\right),\exp\left(-ia\tau\right))$, the differential equation reduces to $\tilde{W}^{\prime}(\tau)=iaM\tilde{W}(\tau)$, with the constant coefficient matrix:
$M=\begin{pmatrix}-\frac{1+S^{2}}{1-S^{2}}&\frac{2S}{1-S^{2}}\\ -\frac{2S}{1-S^{2}}&\frac{1+S^{2}}{1-S^{2}}\end{pmatrix}$. (16.22)
The eigenvalues of $M$ are determined by its trace and determinant:
$\text{tr}M=0,\qquad\det M=-\left(\frac{1+S^{2}}{1-S^{2}}\right)^{2}+\left(\frac{2S}{1-S^{2}}\right)^{2}=-1.$ (16.23)
The eigenvalues of the holonomy generator $iaM$ are $\pm ia$, independent of the boundary parameter $S$. The discretized quark loop consists of $2L$ Dirac matrices, each corresponding to an angular step $\Delta\tau=2\pi/(2L)$. The staggered transformation groups these into $L$ adjacent pairs. The continuous evolution matrix $M$ emerges after the pairing of two adjacent matrices, so it is multiplied $L$ times around the closed loop. The total accumulated parameter $\tau$ evaluates to $L\Delta\tau=\pi$. Consequently, the accumulated phase is twice smaller. The solution over $k$ fundamental torus periods evaluates with the half-angle $\pi ak$:
$\tilde{W}(\pi k)=\cos(\pi ak)I+i\sin(\pi ak)M.$ (16.24)
Transforming back to the original frame with $U(\pi k)$ and computing the trace yields:
$\text{tr}W(\pi k)$
$=\exp\left(i\pi ak\right)\left(\cos(\pi ak)-i\sin(\pi ak)\frac{1+S^{2}}{1-S^{2}}\right)$
$\qquad+\exp\left(-i\pi ak\right)\left(\cos(\pi ak)+i\sin(\pi ak)\frac{1+S^{2}}{1-S^{2}}\right)$
$=2\cos^{2}(\pi ak)+2\sin^{2}(\pi ak)\frac{1+S^{2}}{1-S^{2}}.$ (16.25)
Substituting $S=\sin(\beta)$, the scalar factor reduces to $(1+S^{2})/(1-S^{2})=1+2\tan^{2}(\beta)$. The lower block $\bar{W}(\pi k)$ yields an identical trace. The total trace of the $4\times 4$ spin holonomy evaluates to:
$T_{4}(k)$ $=\text{tr}_{4}\left(\prod\limits_{j}i\gamma_{\mu}v^{\mu}(\tau_{j})\right)$
$=4\left(1+2\sin^{2}(\pi ak)\tan^{2}(\beta)\right).$ (16.26)
16.7 Meson Vertex Operators and Topological Sectors
Taking the logarithm of the oscillating trace $T_{4}(k)$ to include it in the effective action would introduce complex saddle points into the path integral. Instead, treating the bare trace $T_{4}(k)$ as a pre-exponential factor, we expand it analytically into geometric phase factors using the identity $2\sin^{2}(\pi ak)=1-\frac{1}{2}(e^{2\pi iak}+e^{-2\pi iak})$:
$\frac{1}{4}T_{4}(k)=1+\tan^{2}(\beta)-\cos(2\pi ak)\tan^{2}(\beta)$
$=\sec^{2}(\beta)-\frac{1}{2}\tan^{2}(\beta)\left(e^{2\pi iak}+e^{-2\pi iak}\right).$ (16.27)
This expansion resolves the monodromy analytically. The constant coefficients $\sec^{2}(\beta)$ and $\frac{1}{2}\tan^{2}(\beta)$ are real; they dictate the transition amplitudes (residues) of the states but do not shift the complex geometric phase.
To isolate the physical meson states, we insert the corresponding interpolating vertex operator $\Gamma$ ($\gamma_{5}$ for the pseudoscalar pion, $\gamma_{\mu}$ for the vector $\rho$) into the trace:
$T_{\Gamma}(k)=\text{tr}_{4}\left(\Gamma\prod_{j}i\gamma_{\nu}v^{\nu}(\tau_{j})\right).$ (16.28)
The pseudoscalar insertion $\gamma_{5}$ isolates the constant coefficient $\sec^{2}(\beta)$, defining the transition amplitude for the unshifted branch ($q=0$).
The vector insertion $\gamma_{\mu}$ isolates the oscillating phase factors $e^{\pm 2\pi iak}$. When combined with the macroscopic angular momentum phase of the action ($e^{-4\pi iaJk}$), the total phase shift in the winding sum becomes:
$-4\pi aJk\pm 2\pi ak=-4\pi a\left(J\mp\frac{1}{2}\right)k.$ (16.29)
This shifts the target angular momentum in the quantization condition to $u=J+q/2$, where $q\in\{0,-1\}$. This mathematically generates the half-integer spin-orbit splittings for the respective meson Regge trajectories.
16.8 The saddle point equation
Assembling the classical geometric action, we obtain the total monodromy of Minkowski action:
$\Delta S_{\text{geom}}(R,a,\beta)=2\pi ER-\frac{\pi\sigma R^{2}}{a}\left(\beta+\frac{\sin(2\beta)}{2}\right)$
$\qquad-4\pi m_{q}R\cos(\beta)-4\pi aJ-\frac{2a}{3}(\tan\beta-\beta).$ (16.30)
Extremizing this action with respect to $\beta$ isolates the classical boundary condition balancing the string tension against the centrifugal force of the massive endpoints:
$\frac{\partial\Delta S_{\text{geom}}}{\partial\beta}$
$=-\frac{2\pi\sigma R^{2}}{a}\cos^{2}(\beta)+4\pi m_{q}R\sin(\beta)-\frac{2a}{3}\tan^{2}(\beta)=0$
$\implies\sigma R^{2}\cos^{2}(\beta)-2m_{q}aR\sin(\beta)+\frac{a^{2}}{3\pi}\tan^{2}(\beta)=0.$ (16.31)
The bound states of the spinning meson emerge as the meromorphic poles where these geometric series diverge: $\Delta S_{geom}\pm 2\pi aq=2\pi n$.
16.9 Stationary metric and exact spin projection.
The unique property of the exponential solution (16.4) for twistors is the stationary induced metric (16.11), which is strictly independent of the proper time $\tau$. We initially requested only the vanishing variation of the effective action over the temporal period $\Delta\tau=2\pi$, but the exact solution we have found is, in fact, strictly stationary. Not only does its monodromy vanish, but the metric itself is entirely independent of the proper time $\tau$.
Moreover, the branch point degree $a$ (related to the macroscopic target-space rotation angle by $\alpha=4\pi a$) was left completely arbitrary by the saddle point equations for the remaining moduli. Previously, we considered this continuous degeneracy as a flat valley required to dynamically generate a resonance pole in the steepest descent integral, which would necessitate conjecturing that all higher functional derivatives vanish. Now, we can finalize that line of thought.
To extract a physical meson state of definite angular momentum $J$, we introduced the Fourier projector (15.1) of our twisted amplitude $A(\alpha)$ onto the given $J$ state:
$A_{J}=\int_{0}^{2\pi}d\alpha\,\sum_{k}e^{-\mathrm{i}k\alpha J}A_{k}(\alpha)$
$=\sum_{k=1}^{\infty}\int_{0}^{2\pi}d\alpha\exp\left(\mathrm{i}k\alpha\Phi_{\star}(E,J)/(4\pi)\right)$
$\propto\frac{4\pi\mathrm{i}}{\Phi_{\star}(E,J)}\sum_{k=1}^{\infty}\frac{1-\exp\left(\mathrm{i}k\Phi_{\star}(E,J)/2\right)}{k}$ (16.32)
Here $\Phi_{\star}(E,J)$ is the value of $\Phi(E,J,\mathcal{K},\beta)$ at its extremum as described by above saddle point equations for moduli $\mathcal{K},\beta$. The last sum logarithmically diverges at large windings $k$, which justifies our WKB approximation used in estimating the Hessian and neglecting deviations from this classical trajectory. The coefficient in front of this logarithmically divergent factor is
$A_{J}\propto\frac{4\pi\mathrm{i}\log k_{\text{max}}}{\Phi_{\star}(E,J)}$ (16.33)
Remark 16.1.
Hessian estimate and Quantum Fluctuations
A more accurate estimate of the divergent sum over windings must account for the transverse quantum fluctuations of the trajectory around this flat classical valley. Because the classical orbit length scales linearly with the winding number $k$, the stability matrix (the Hessian $H$) is directly proportional to $k$. We presume that the fluctuating part of the twistor fields is strictly periodic, so that it will not contribute to the monodromy The fluctuating part of the action will, as usual , start with quadratic terms (the linear terms vanishing because we minimize over parameters of the classical solutions we have considered above). Therefore, the $k$-fold integral over the period $\Delta\tau=2\pi$ for this fluctuative part will be independent of $a$ and linear in $k$. Let us elaborate. The effective action for the fluctuative part will no longer be constant in $\tau$, but it will be periodic, which will result in linear $k$- dependence of the integral over multiple periods $0<\tau<2\pi k$.
The scaling factor for the integration measure can be extracted via the $\zeta$-regularization of the functional determinant of this Hessian. Using the canonical value $\zeta(0)=-1/2$ for the trace over the 1D loop boundary modes, this yields the fluctuation weight: